OpenStudy (calculusxy):

Simplify the following expression: 9^4 X 6^3 ------------ 9^3 X 6

OpenStudy (calculusxy):

Yes

OpenStudy (calculusxy):

It's really dividing.

OpenStudy (calculusxy):

@agent0smith

OpenStudy (agent0smith):

Simplify it using exponent rules.

OpenStudy (calculusxy):

This is a fraction, it comes with only one answer. The answer choices are: 54, 486, 1,816, and 324.

OpenStudy (anonymous):

9^3x9X6^3/9^3X6 == 9X6^2 = 9x36

OpenStudy (calculusxy):

There was this rule that you are supposed to subtract the exponents when there is the same base. But I don't know how to appropriately apply that rule.

OpenStudy (agent0smith):

Try it.

OpenStudy (calculusxy):

So since both 9^4 and 9^3 have 9, I did the subtraction and got 9^1. The problem comes to subtracting 6^3 and 6. How do I do that?

OpenStudy (anonymous):

6 is 6^1

OpenStudy (agent0smith):

9^1 is the same as...? 6 has an exponent of..?

OpenStudy (anonymous):

its 324, sorry i thought it was 2 different problems

OpenStudy (calculusxy):

9^1=9 but I don't get what you mean by "6 has an exponent of...?"

OpenStudy (agent0smith):

Well if 9^1 = 9, then what must be the exponent on 6?

OpenStudy (calculusxy):

I am not sure about that. My best answer was 6^3. But I am not sure whether I am correct or not

OpenStudy (calculusxy):

@agent0smith Are you here?

OpenStudy (agent0smith):

Well if 9^1 = 9, then what must be the exponent on 6? ^^^^Look

OpenStudy (calculusxy):

Finally got it! It was supposed to be 6^1=6 so I subtracted 3-1=2. So I multiplied 9*36=324

OpenStudy (agent0smith):

Yep.

OpenStudy (calculusxy):

Thanks for your help.

OpenStudy (agent0smith):

^it's more important he understands how to do it, than someone giving the right answer